The number of chocolate chips in an​ 18-ounce bag of chocolate chip cookies is approximately normally distributed with a mean of 1252 chips and standard deviation 129 chips. ​(a) What is the probability that a randomly selected bag contains between 1100 and 1500 chocolate​ chips, inclusive? ​(b) What is the probability that a randomly selected bag contains fewer than 1025 chocolate​ chips? ​(c) What proportion of bags contains more than 1200 chocolate​ chips? ​(d) What is the percentile rank of a bag that contains 1425 chocolate​ chips? ​(a) The probability that a randomly selected bag contains between 1100 and 1500 chocolate​ chips, inclusive, is ______?
Answer #1

solution Given M=1252 G 129 a Plnoo xxxgood. p/ 1100-1252 xall, 1500-12521 P(-1.18 <Z4 1.92) (z= x-ll) = P(Z <1-92)-P(Z <-1.1please like if it helps please

Thank you